FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF ARGON-IMIDAZOLE

\underline{RYAN WELCH} and HELEN O. LEUNG, Department of Chemistry, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000.

Guided by the spectroscopic constants obtained in a millimeter-wave study,^{a} the rotational spectrum of Ar-imidazole has been recorded in the 8 - 18 GHz region using a pulsed molecular beam, Fourier transform microwave spectrometer. The nuclear quadrupole hyperfine structure due to the quadrupolar ^{14}N nuclei has been resolved for these low J transitions. The nuclear quadrupole coupling constants for each nitrogen nucleus in the complex will be compared with those in the imidazole monomer to derive molecular properties of the complex.

aW. Caminati, S. Melandri, A. Millemaggi, and P. G. Favero, Chem. Phys. Lett. 294, 377 (1998).